Assumptions                                     Qd = 10000/P      &nbs

Assumptions

                                    Qd = 10000/P                         Qs = 100P

  1. What is Equilibrium Q and P

  1. If a 25% sales tax is placed on the Good. What is new Pc, Pp, and Q. Who pays more of the tax?
Answer #1

a)

Set Qd=Qs for equilibrium

10000/P=100P

100=P2

P=$10 (equilibrium price)

Qd=10000/P=10000/10=1000

Qs=100P=100*10=1000

So, Equilibrium quantity=Qd=Qs=1000

b)

Sales tax is paid by the consumers.

P will be replaced by P*(1+25%) i.e. 1.25P

Qd'=10000/1.25P=8000/P

Set Qd'=Qs for equilibrium

8000/P=100P

80=P2

P=8.944272 or say $8.94

Qd'=8000/P=8000/8.944272=894.42 units

Qs=100P=100*8.944272=894.42 units

New equilibrium quantity =Q=894.42

Price paid by consumers=Pc=P*(1+25%)=8.94*1.25=$11.18

Price received by sellers=Pp=P=$8.94

Tax burden on buyers=11.18-10=$1.18 per unit

Tax burden on sellers= 10-8.94=$1.06

We can say that burden of tax is more on buyers. So, buyers pay more of tax in this case.